Find the period of the valued function satisfying
f(x)+f(x+4)=f(x+2)+f(x+6)

Grade:

1 Answers

Arun Kumar IIT Delhi
9 years ago
Hello Student,
f(x)+f(x+4)=f(x+2)+f(x+6)
=>f(x+2)+f(x+6)=f(x+4)+f(x+8)
adding both
we get
f(x)=f(x+8)
Thanks & Regards
Arun Kumar
Btech, IIT Delhi
Askiitians Faculty